OpenStudy (anonymous):

Perfom the indicated operation. Simplify as possible. 8x/7x+28 - x-1/x+4 Am completely stuck

OpenStudy (anonymous):

This right? \[\frac{ 8x }{ 7x+28 } - \frac{ x-1 }{ x+4 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First you need a lowest common denominator.

OpenStudy (anonymous):

i got 8x/7

OpenStudy (anonymous):

x+4 can multiply into 7x+28, so 7x+28 is the LCD

OpenStudy (anonymous):

So you need to multiply the right side by 7 to get the same denominator.

OpenStudy (anonymous):

\[\frac{ x-1 }{ x+4 } x \frac{ 7 }{ 7 } = \frac{ 7x-7 }{ 7x+28 }\]

OpenStudy (anonymous):

So now you have \[\frac{ 8x }{ 7x+28 } - \frac{ 7x-7 }{ 7x+28 }\]

OpenStudy (anonymous):

so is the answer x-7/7x+28?

OpenStudy (anonymous):

Correct, just subtract across the top.

OpenStudy (anonymous):

so that is the final answer?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Remember when subtracting fractions, you need to get an LCD first. Then the problem is easy.

OpenStudy (anonymous):

okay i really need help on this one

OpenStudy (anonymous):

11/x+3 = 19x+18/x^2+7x+12 - 18/x+4